HETAV PATEL

Grade 11,

Water is flowing in a river at 2.5m/s. The river is 40m wide with an average depth of 6m. compute the power available from the water current.

Water is flowing in a river at 2.5m/s. The river is 40m wide with an average depth of 6m. compute the power available from the water current.
 

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\$consider a section dx of the water$ \\mass=40*6*dx*\rho \\weight=mg \\$height of center of mass$=h/2=3 \\potential energy stored=mgh/2 \\Power=dP/dt=40*6*{dx \over dt}*\rho g*3 \\Power=6*10^6($considering g=10,\rho=1000kg/m^3$) \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty